Fill in the blanks.
 
(c) 12 + (–12) = ____
(d) (–4) + ____ = –12
(e) ____ – 15 = – 10
 
iaemfuewgnuersiojmhnbiorjhmneru

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



(c) 12 + (-12) = 0

(d) (-4) + (-8) = -12

(e) 5 – 15 = -10

​Thanks